K, = dissociation constant kon R-receptor L ligand off RIIL K, Total Rec:R,-[R]+[RL] = - Kd ILI QUESTIONS Youve identified a drug that has the potential to bind and inhibit 5 different histamine receptors found on human mast cells. The drug if ingested orally typically reaches a concentration of 1 x 108 Molar in the blood stream. Given the K of the drug for each of the 5 receptors listed below, calculate the percentage of receptors occupied with a drug concentration of 1 X 10-8 Molar Percentage of Receptors Occupied (with 108 Molar drug concentration) RECEPTOR Receptor A K, 1 micromolar Receptor B K 10 micromolar Receptor C K, 100 micromolar Receptor D K1 nanomolar Receptor E K, 10 nanomolar
Answer #1

In the question, it is mentioned that ([RL]/RT)= [ 1 / {1+(Kd / [L])} ]

([RL]/RT) represents the fraction of receptors occupied with a drug.

So,

The fraction of receptors occupied with a drug= [ 1 / {1+(Kd / [L])} ]= [1 / {([L] + Kd) / [L]}] = {[L] / ([L]+Kd)} Where, [L]= Drug concentration, Kd= Dissociation constant.

To solve these problems, we can use the above-mentioned formula.

Calculation for receptor A:

[L]= 1\times10-8 Molar= 1\times10-8\times106 micromolar [As 1 Molar= 106 micromolar] = 10-2 micromolar.

Kd= 1 micromolar

So, the percentage of receptor A occupied with the drug= {[L] / ([L]+Kd)}= {10-2/(10-2+1)}

= 9.9\times10-3= (9.9\times10-3\times100) %= 0.99 % [Answer]

Therefore, 0.99% of receptor A is occupied with the drug.

Calculation for receptor B:

[L]= 1\times10-8 Molar= 1\times10-8\times106 micromolar [As 1 Molar= 106 micromolar] = 10-2 micromolar.

Kd= 10 micromolar

So, the percentage of receptor B occupied with the drug= {[L] / ([L]+Kd)}= {10-2/(10-2+10)}

=9.9\times10-4= (9.9\times10-4\times100) %= 0.099 % [Answer]

Therefore, 0.099% of receptor B is occupied with the drug.

Calculation for receptor C:

[L]= 1\times10-8 Molar= 1\times10-8\times106 micromolar [As 1 Molar= 106 micromolar] = 10-2 micromolar.

Kd= 100 micromolar

So, the percentage of receptor C occupied with the drug= {[L] / ([L]+Kd)}= {10-2/(10-2+100)}

= 9.9\times10-5= (9.9\times10-5\times100) %= 0.0099 % [Answer]

Therefore, 0.0099% of receptor C is occupied with the drug.

Calculation for receptor D:

[L]= 1\times10-8 Molar= 1\times10-8\times109 nanomolar [As 1 Molar= 109 nanomolar] = 10 nanomolar.

Kd= 1 nanomolar

So, the percentage of receptor D occupied with the drug= {[L] / ([L]+Kd)}= {10/(10+1)}= 0.91 =(0.91\times100) %= 91% [Answer]

Therefore, 91% of receptor D is occupied with the drug.

Calculation for receptor E:

[L]= 1\times10-8 Molar= 1\times10-8\times109 nanomolar [As 1 Molar= 109 nanomolar] = 10 nanomolar.

Kd= 10 nanomolar

So, the percentage of receptor E occupied with the drug= {[L] / ([L]+Kd)}= {10/(10+10)}= 0.5 =(0.5\times100) %= 50% [Answer]

Therefore, 50% of receptor E is occupied with the drug.